OpenStudy (anonymous):

cos2A. cos2B + sin^2 (A-B) - sin^2 (A+B) is =?

OpenStudy (anonymous):

and could u plz give us the answer to check it?

OpenStudy (anonymous):

the answer should be cos(2A + 2B)

OpenStudy (experimentx):

sin(a-b) - sin (a+b) = -2 cos a sin b

OpenStudy (apoorvk):

yeah, you made a slight 'sign' error @mukushla

OpenStudy (experimentx):

sin^2 (A-B) - sin^2 (A+B)= -sin2A cos2A

OpenStudy (anonymous):

yes thats right try to get sin^2 (A-B) - sin^2 (A+B)=-sin2A cos2A with expanding it then u have cos2A. cos2B-sin2A cos2A=cos(2(A+B))

OpenStudy (anonymous):

plzzcan u upload the full steps

OpenStudy (anonymous):

sin^2 (A-B) - sin^2 (A+B)=-sin2A cos2A??????????

OpenStudy (experimentx):

just work out yourself .. this is simple a^2-b^2 algebra

OpenStudy (anonymous):

hw?

OpenStudy (anonymous):

(a-b)(a+b)

OpenStudy (experimentx):

sin^2 (A-B) - sin^2 (A+B) = (sin (A-B) - sin(A+B))(sin (A-B) + sin (A+B))

OpenStudy (anonymous):

got itthxxx